OpenStudy (anonymous):

Find the square root of (999993)(999991)(999989)(999987)+16.

OpenStudy (lgbasallote):

have you tried a device called "calculator"? unless you're interested to learn how to solve that manually

OpenStudy (anonymous):

yes, there is.....

OpenStudy (lgbasallote):

well then, the first step is to solve (999993)(999991) manually

OpenStudy (anonymous):

Write 1mm as x then (x-7)(x-9)(x-11)(x-13)+16 is (x^2-20x+95)^2

OpenStudy (anonymous):

well, that's right....

OpenStudy (anonymous):

999980000095

OpenStudy (anonymous):

If we let y = 999990, the expression boils down to \[\sqrt{(y+3)(y+1)(y-1)(y-3) +16}\]\[=\sqrt{(y^2-1)(y ^{2}-9)+16}\]\[=\sqrt{y^4 - 10y^2+9+16}\]\[=\sqrt{y^4-10y^2+25}\]\[=y^2-5\]then,\[=(999990)^{2}-5=988880000100-5=988880000095\]